How many grams of water should be produced from 2 g H2 and 16 g O2?
1
Expert's answer
2020-10-26T14:49:28-0400

2H2+O2=2H2O

At first, the limiting reagent should be found.

n=m/M,

n(H2)=2/2=1(mol),

n(O2)=16/32=0.5(mol).

According to the chemical equation, the ratio of moles of H2:O2 should be 2:1 (or 1:0.5), the same as it is given. So, all the reagents are consumed.

m(H2O)=2+16=18(g).